Fact Pattern:

Zeta Company is preparing its annual profit plan. As part of its analysis of the profitability of individual products, the controller estimates the amount of overhead that should be allocated to the individual product lines from the information given in the next column:

Wall

Specialty

Mirrors

Windows

Units produced

25

25

Material moves per product line

5

15

Direct labor hours per unit

200

200

Budgeted materials handling costs

$50,000

Under activity-based costing (ABC), Zeta’s materials handling costs allocated to one unit of wall mirrors would be

$2,500

$500

$1,000

$1,500
Answer #1

Material handling cost per movement = 50000 /(25*5+25*15) = $100

Material handling cost allocated to one unit of wall mirror = 100*5 = $500